OpenStudy (anonymous):

MEDALLLLL!!! 1There are 100 runners entered in a marathon. How many different groups of three runners can finish in first, second, and third? Does this problem involve permutations or combinations? 2There are 100 runners entered in a marathon. Trophies are awarded for first place, second place, and third place. In how many different ways could the runners receive trophies? Does this problem involve permutations or combinations?

OpenStudy (anonymous):

The question states "groups of 3". In other words A first, B second and C third is exactly the same group of three as B first C second and A third. Where order of selection doesn't matter you use combinations. Where order of selection does matter you use permutations. Therefore the answer = 100C3 or choose any three from 100 = 100! / (3!97!) = (100 * 99 * 98 * 97!) / (3 *2 * 1 *97!) = 100* 99 * 98 /6 = 161700

OpenStudy (anonymous):

You would do the second part the same way

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

you would do it the same way

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

:)